OpenStudy (anonymous):

Will fan and Medal! I have no idea how to solve this it confuses me.. The sun produces 3.9 ⋅ 1033 ergs of radiant energy per second. How many ergs of radiant energy does the sun produce in 1.55 ⋅ 107 seconds?

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

is it \[3.9\times 10^{33}\]?

OpenStudy (misty1212):

and \(1.55\times 10^7\)?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

this may be confusing, so lets make it simple ok?

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

if you drive 60 miles an hour, how far do you go in two hours?

OpenStudy (anonymous):

120?

OpenStudy (misty1212):

yes , and you got it by multiplying

OpenStudy (anonymous):

yes i did

OpenStudy (misty1212):

same idea here, you need to multiply these numbers

OpenStudy (anonymous):

4028.7

OpenStudy (misty1212):

\[3.9\times 10^{33}\times 1.55\times 10^7\]

OpenStudy (anonymous):

165.85

OpenStudy (misty1212):

i really doubt it is the answer you got

OpenStudy (misty1212):

how to you multiply these? you do it by multiplyiing \[3.9\times 1.55\] then adding the exponents on the powers of ten

OpenStudy (misty1212):

\[3.9\times 1.55\times 10^{33+7}\] \[3.9\times 1.55\times 10^{40}\]

OpenStudy (misty1212):

i would use a calculator for the first part

OpenStudy (anonymous):

6.045

OpenStudy (misty1212):

right final answer is \[\huge 6.045\times 10^{40}\]

OpenStudy (anonymous):

hay! that wasn't so hard! You're awesome! thanks so much :)

OpenStudy (misty1212):

no it is not so hard when you know what to do your are welcome

OpenStudy (anonymous):

i have one last problem

OpenStudy (anonymous):

Find the value of the following expression: (28 ⋅ 3−5 ⋅ 60)−2 ⋅ 3 to the power of negative 2 over 2 to the power of 3, whole to the power of 4 ⋅ 228

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

can you post a screen shot or something?

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

lol

OpenStudy (anonymous):

Write your answer in simplified form. Show all of your steps

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

\[\huge \left(\frac{3^{-2}}{2^3}\right)^4\]

OpenStudy (anonymous):

yes and the rest is (28 ⋅ 3−5 ⋅ 60)−2 ⋅ ⋅ 228

OpenStudy (misty1212):

i can't read that, but for the one above if you multiply the exponents you get \[\frac{3^{-8}}{2^{12}}\]

OpenStudy (anonymous):